OpenStudy (prettyboy14):

A car is driving at a speed of 50mi/h. What is the of the car in feet per minute?

OpenStudy (anonymous):

How many feet are in one mile? How many minutes are in one hour?

OpenStudy (anonymous):

5280 feet in a mile

OpenStudy (anonymous):

60 minutes in a hour

OpenStudy (anonymous):

\[50 \frac{ mi }{ h } * \frac{ 5280 ft }{ 1 mi } * \frac{ 1 h }{ 60 \min }\]

OpenStudy (anonymous):

60 mph = 88 ft/sec \[\frac{50}{60}*88*60=4400 \text{ ft}\]